# Parameterized Prompt for Incremental Object Detection

### v2

@@ 4.3 Parameterized Prompt Fusion for Incremental Learning
The parameterized prompt also faces catastrophic forgetting during incremental learning. To address the forgetting, we introduce model fusion after each task training. During the fusion process, we aim to retain the important parameters of each task and average the consistent parameters across tasks. Furthermore, we introduce a sparse loss to concentrate the knowledge of each task in important parameters for facilitate the model fusion.
@@
During the $T_{t}$ fusion process, we preserve critical parameters guided by $\mu_{t}$ and $\mu_{t-1}^{f}$, and average consistent parameters based on $\gamma_{t}$ and $\gamma_{t-1}^{f}$. To preserve critical parameters, we first sort the values in $\mu_{t-1}^{f}$ and select the ${\text{top-}}k$% indices, denoted as $\mathcal{I}_{t-1}^{f}$. The corresponding parameter in $\theta_{t-1}^{f}$ are preserved with priority at $\mathcal{I}_{t-1}^{f}$. Next, we sort $\mu_{t}$ and identify the ${\text{top-}}l$% indices, denoted as $\mathcal{I}_{t}$. The parameter in $\theta_{t}$} are preserved at $\mathcal{I}_{t}$, excluding any overlap with $\mathcal{I}_{t-1}^{f}$. To average consistent parameters, we locate positions where $\gamma_{t}=\gamma_{t-1}^{f}$, indicating directional consistency. At these positions, excluding those already reserved for preservation (i.e., $\mathcal{I}_{t-1}^{f}\cup\mathcal{I}_{t}$), we take the average of $\theta_{t}$ and $\theta_{t-1}^{f}$. Finally, all remaining undecided parameters are assigned the corresponding values from $\theta_{t-1}^{f}$. The overall fusion process can be formally expressed as follows:
@@
where $i$ denotes the $i$-th parameter in the parameterized prompts. The pseudo-code for parameterized prompt fusion is outlined in Appendix D.1.
@@
where $\lambda$ is the hyperparameter controlling the sparsity level, and $\theta_{j}$} refers to the parameterized prompts in $j^{{\text{th}}}$} decoder layer.
4.4 Pseudo Labeling for Mining co-occurring Objects
In IOD’s co-occurring scenarios, unlabeled previous task objects may appear in the background of current task images. Properly mining these previous task objects can significantly reduce forgetting, while treating these objects as background can lead to more severe forgetting. We propose a simple heuristic method to mine the knowledge of co-occurring objects in the background.
In $T_{t}$, we employ the detector to infer on each training sample, obtaining predictions: ${{\hat{y}}_{i}}=\left\{{{{\hat{s}}_{i}},{{\hat{b}}_{i}}}\right\}$. Here, ${\hat{s}}_{i}$}} represents the score for the highest-scoring category, and ${\hat{b}}_{i}$}} provides the bounding box coordinates for this prediction. pseudo-labeling mechanism [5, 9, 17] sets a threshold $\tau$ to filter predictions with ${\hat{s}}_{i}$}} higher than this threshold as pseudo label ${{\tilde{y}}_{i}}=\left\{{{{\tilde{c}}_{i}},{{\tilde{b}}_{i}}}\right\}$. The threshold $\tau$ ensures that only the most reliable predictions are used when generating pseudo labels. Here, ${\tilde{c}}_{i}$}} represents the pseudo label’s category name, ${\tilde{b}}_{i}$}} is the bounding box coordinates for this pseudo label. In task $T_{t}$}}, pseudo labels incorporate the knowledge from previous tasks $\left\{{{T_{1}}\ldots{T_{t-1}}}\right\}$, effectively alleviating the detector’s forgetting.
Table 1: Average precision ($A{P_{50}}$, %) is compared on the PASCAL VOC2007 dataset under single-step settings of 19+1, 15+1, 10+10, and 5+15. Moreover, we add the superscript to the accuracy that may be overestimated. The reasons for the overestimation are detailed in the eval metrics.
@@ 5.1 Experimental Settings
Eval metrics. Followed by [3], we use the mean average precision at an IOU threshold of 0.5 ($A{P_{50}}$, %) as the metric. For PASCAL VOC2007, following previous works [13], we provide the $A{P_{50}}$ of the current task classes and the previous task classes to better reflect the method’s stability and plasticity. There are two evaluation methods for obtaining task precision: validating on the entire test set versus using task-specific test subsets. The second method yields higher precision for the same model. We adopt the first method and add superscript to results from the second method to ensure fair comparison. For MS COCO, following previous works [12], we provide the $A{P_{50}}$ of all trained classes after training at each stage.
Implementation details. We implement our proposed method based on Deformable-DETR [34] pre-trained on the MS COCO dataset and Co-DETR [36] pre-trained on the Objects365 dataset [22], both obtained from HuggingFace. The large-scale Objects365 dataset contains 365 object categories and over 600,000 images, making it a suitable pre-training source for incremental learning experiments on MS COCO dataset. Furthermore, since the official MD-DETR implementation is only available on Deformable-DETR, we port MD-DETR [3] on Co-DETR for a fair comparison.
@@ 5.2 Comparison
Table 4: Average precision ($A{P_{50}}$, %) is compared on the PASCAL VOC2007 dataset under multi-step settings of 10+5+5 and 5+5+5+5. We add the superscript to the accuracy that may be overestimated. The reasons for the overestimation are detailed in the eval metrics.
@@
Single-step setting. We compare three single-step scenarios on the PASCAL VOC2007 dataset, where the co-occurrence levels gradually increase in the 19+1, 15+5, and 10+10 settings. As shown in Tab. 1, P2IOD with MS COCO and Objects365 pretrained detectors achieve excellent performance across all experimental settings. Compared to the prompt-based MD-DETR, P2IOD achieves accuracy improvements of 1.6% / %, 2.5% / 3.9%, and 8.0% / 5.2% in the respective scenarios, indicating that the performance advantage of P2IOD becomes increasingly significant as the co-occurrence level rises. This trend further demonstrates that P2IOD mitigates the interference caused by prompt-pool confusion in co-occurring scenarios.
Multi-step setting. We compare the multi-step settings on PASCAL VOC2007 and MS COCO datasets. Tab. 4 shows that on PASCAL VOC2007, the performance degradation of MD-DETR becomes increasingly severe as the number of incremental steps grows. The degradation stems from prompt pool confusion, which intensifies as the number of tasks increases. In contrast, our method effectively mitigates such confusion, consistently achieving superior performance across all settings. For the MS COCO dataset, as shown in Tab. 2, P2IOD also exhibits the aforementioned advantages. Moreover, P2IOD consistently outperforms other existing approaches across different settings on both datasets, demonstrating its effectiveness and the strong potential of prompt-based techniques in IOD.
@@ 5.3 Analysis
Ablation Study. In Tab. 3, categories 1-5 reflect the stability, while categories 6-20 mainly reflect plasticity. After introducing the pseudo-labeling method (b), due to the lack of learnable parameters, pseudo-labeling not only fails to enhance stability but also interferes with current task learning, reducing plasticity. The parameterized prompt structure (c) increases the accuracy of categories 6-20 by 9.5%, significantly enhancing plasticity, but the accuracy of categories 1-5 drops by 2.6%, indicating that forgetting still exists.Model fusion (d) alleviates the forgetting problem and balances stability and plasticity to some extent. Sparse loss (f) compresses knowledge into important parameters, and its combination improves overall task accuracy by 8.9% compared to the baseline. Furthermore, removing the pseudo-labeling (e) results in a significant decrease in stability. The observation demonstrates that our method, by alleviating prompt-pool confusion, allows the pseudo labeling mechanism to effectively mine old-class objects in the background without introducing adverse effects.
Impact of Hidden Layer Dimension. We analyze the impact of the hidden layer dimension in the parameterized prompt structure. The dimension of the hidden layer influences the degree of dimensionality reduction applied to the proposal. We conduct experiments in the PASCAL VOC2007 under the 5+5+5+5 setting. In Fig. 4, as the hidden layer dimension increases, the model’s accuracy initially improves and then declines, suggesting that a moderate increase in the hidden dimension helps retain critical information, while an overly large dimension introduces redundant information that interferes with prompt generation. As the hidden dimension increases, the number of parameters in the parameterized prompt increases accordingly. Our method achieves a favorable trade-off between performance and parameter efficiency (76.3%, 1.1M) when the hidden dimension is set to 64. In contrast, MD-DETR [3] requires more parameters, while simultaneously achieving lower accuracy (61.5%, 1.8M).
@@
Cross-Task Prompt Comparison. We compare the distribution similarity of prompts across tasks between our method and MD-DETR. We conduct this experiment on PASCAL VOC2007 (5+5+5+5) and use Maximum Mean Discrepancy (MMD) [8] to evaluate the distribution similarity of prompts, with the average MMD (A-MMD) across all tasks as the evaluation metric. A larger A-MMD value indicates a more significant prompt diversity. As shown in Fig. 5, the diversity of prompt distributions in our P2IOD increases with the depth of decoder layers, and the diversity at the final layer is significantly higher than that of MD-DETR. Our method generates category-independent prompts in shallow layers and category-related prompts in deep layers, aligning well with the multi-layer decoder architecture, while MD-DETR is constrained by its pool structure and struggles to match this characteristic. Furthermore, our method can generate more diverse prompts at the final layer used for object prediction. The variations in prompt distributions highlight the effectiveness of our approach.
@@ 5.4 Visualized Comparison
We analyze the visualized comparison between P2IOD and MD-DETR to illustrate that the confusion issue is being addressed. Specifically, in Fig. 3, the visualizations of MD-DETR exhibit numerous false positives, indicating that the confused prompts pool introduces incorrect prompts into the detector, thereby increasing scores for irrelevant categories. In contrast, P2IOD significantly reduces such false positives, demonstrating the effectiveness of our approach in mitigating confusion. We also observe that although P2IOD and MD-DETR have nearly identical confidence for detecting people in the first task, as the tasks increase, the confidence in MD-DETR rapidly declines, indicating that the confusion in MD-DETR affects the confidence of positive targets. In contrast, the confidence in P2IOD remains almost unchanged, proving that our method is unaffected by confusion.
@@ 6 Conclusion
In this study, we identify a severe confusion issue within the prompts pool under co-occurring scenarios, which exacerbates catastrophic forgetting as the degree of co-occurrence and the number of learning steps increase. To address this issue, we argue that prompts in IOD should not be assigned to individual tasks exclusively but should exhibit adaptive consolidation properties across tasks, with constrained updates. We propose a parameterized prompt structure and parameterized prompt fusion to validate our hypothesis. Experiments on multiple datasets demonstrate that our framework exhibits superior performance compared to state-of-the-art methods. To our knowledge, this is the first work addressing prompts pool confusion in incremental object detection, laying a foundation for broader prompt-based IOD applications.
@@
7 Implementation Details
The pre-trained detectors use the same original hyperparameter settings as those in their respective papers [34, 36]. We employ the Adam optimizer with a weight decay of 0.0001. The learning rate is set to 0.0001 for the class embeddings and the parametrized prompts, and a lower learning rate of 0.00001 is used for the bounding box embeddings, while freezing the remaining parameters. For training Deformable-DETR [34] on the PASCAL VOC2007 dataset [6], we train each incremental task for 100 epochs, with the learning rate dropped to 0.1 of its original value at the 80th epoch. For training Co-DETR [36] on the PASCAL VOC2007 dataset, we train each incremental task for 12 epochs, with the learning rate dropped to 0.1 of its original value at the 7th epoch. For training Co-DETR on the MS COCO dataset, we train each incremental task for 1 epoch. It is worth noting that in the 19+1 setting on PASCAL VOC2007, the second task contains only 279 images. Due to the limited data, Deformable-DETR tends to overfit during training. To mitigate this issue, we set the hyperparameters of the focal loss to $\alpha=0.5$ and $\gamma=3.0$. The reported accuracy is the average of three independent trials. All experiments are conducted using an NVIDIA RTX 4090 GPU with a batch size of 32.
On the PASCAL VOC2007 dataset [6], we train each incremental task for 100 epochs, with the learning rate drop at the 80th epoch. On the MS COCO dataset [15], we train each incremental task for 12 epochs, with the learning rate drop at the 10th epoch. For the comparison experiment in PASCAL VOC2007 dataset 19+1, 15+5, and 10+10 settings, we adopted the accuracy as reported in the MD-DETR. But due to the lack of accuracy reported in other scenarios, we reproduce MD-DETR and conduct experiments under our experimental settings for the PASCAL VOC2007 dataset 5+5+5+5 and 10+5+5 settings, as well as the MS COCO dataset 40+20+20 and 40+10+10+10+10 settings. To ensure a fair comparison, we set the length of prompts to be consistent with the setting in MD-DETR. The settings for pseudo-labeling method follow those established in MD-DETR, with $\tau$ set to 0.65. Additionally, in the 19+1 setting on PASCAL VOC2007, the second task contains only 279 images. The limited data leads to overfitting during training with Deformable-DETR. To mitigate this issue, we set the hyperparameters of the focal loss to $\alpha=0.5$ and $\gamma=3.0$. The reported accuracy is the average over three independent trials. All experiments were conducted using an NVIDIA RTX 4090 GPU with a batch size of 32.
@@ 7 Implementation Details
8 Additional Experiment Results
8.1 Single-step Comparison on MS COCO dataset
We compare two single-step scenarios on the MS COCO dataset, namely the 40+40 and 70+10 settings. As shown in Table 5, P2IOD performs excellently across all experimental settings. Compared to the prompt-based MD-DETR, P2IOD achieves $A{P_{50}}$ accuracy improvements of 5.4% and 6.3% in the two scenarios, demonstrating that the proposed method effectively addresses prompt-pool confusion. Our method achieves a maximum $A{P_{50}}$ of 71.1% and 71.9% in the two settings, outperforming other methods by 8.2% and 8.0%. This proves the great potential of prompt-based methods in IOD.
8.2 Impact of Parameterized Prompt Fusion Threshold.
@@ 8.1 Single-step Comparison on MS COCO dataset
We quantitatively analyze the impact of the ${\text{top-}}k$ and ${\text{top-}}l$ in parameterized prompt fusion. The threshold determines the proportion of the previous and current parameterized prompt structures retained during fusion. To clarify the impact of one-step fusion on accuracy, we perform experiments in the PASCAL VOC2007 under 10+10 setting. As shown in Tab. 6, we observe that retaining only the current task’s parameters (no fusion) or previous task’s parameters (${\text{top-}}k=1.0$) lead to degraded performance. When ${\text{top-}}k=0.0$ and ${\text{top-}}l=0.0$, the method averages the consistent parameters, improving both the current and previous tasks compared to the non-fused approach, demonstrating that averaging consistent parameters enhances generalization. Furthermore, by comparing different values of ${\text{top-}}k$ and ${\text{top-}}l$, we find that increasing ${\text{top-}}k$ and ${\text{top-}}l$ to a certain extent further improved performance. The accuracy improvement indicates that retaining key parameters from each task helps preserve task-specific knowledge. We observe the best performance at ${\text{top-}}k=0.7$ and ${\text{top-}}l=0.3$, resulting in a 1.08% accuracy increase over the non-fused method.
8.3 Impact of $\lambda$ in Sparse Loss.

### v4

@@ 4.3 Parameterized Prompt Fusion for Incremental Learning
The parameterized prompt also faces catastrophic forgetting during incremental learning. To address the forgetting, we introduce model fusion after each incremental training process. During the fusion process, we aim to retain the important parameters of each task and average the consistent parameters across tasks. We also introduce a sparse loss to concentrate the knowledge of each task in different parameter subsets to facilitate the model fusion.
@@
During the $T_{t}$ fusion process, we preserve critical parameters guided by $\mu_{t}$ and $\mu_{t-1}^{f}$, and average consistent parameters based on $\gamma_{t}$ and $\gamma_{t-1}^{f}$. To preserve critical parameters, we first sort the values in $\mu_{t-1}^{f}$ and select the ${\text{top-}}k$% indices, denoted as $\mathcal{I}_{t-1}^{f}$. The corresponding parameter in $\theta_{t-1}^{f}$ are preserved with priority at the indices in $\mathcal{I}_{t-1}^{f}$. Next, we sort $\mu_{t}$ and identify the ${\text{top-}}l$% indices, denoted as $\mathcal{I}_{t}$. The parameter in $\theta_{t}$} are preserved at the indices in $\mathcal{I}_{t}$, excluding any overlap with the indices in $\mathcal{I}_{t-1}^{f}$. To average consistent parameters, we locate positions where $\gamma_{t}=\gamma_{t-1}^{f}$, indicating directional consistency. At these positions, excluding those already reserved for preservation (i.e., $\mathcal{I}_{t-1}^{f}\cup\mathcal{I}_{t}$), we take the average of $\theta_{t}$ and $\theta_{t-1}^{f}$. Finally, all remaining undecided parameters are assigned the corresponding values from $\theta_{t-1}^{f}$. The overall fusion process can be formally expressed as follows:
@@
where $i$ denotes the $i$-th parameter in the parameterized prompts. The pseudo-code for parameterized prompt fusion is outlined in Appendix C.1.
@@
where $\lambda$ controlls the sparsity level, and $\theta_{j}$} refers to the parameterized prompts in the $j^{{\text{th}}}$} decoder layer.
Table 1: Average precision ($A{P_{50}}$, %) is compared on the PASCAL VOC2007 dataset under single-step settings of 19+1, 15+1, 10+10, and 5+15. We add the superscript to the accuracy that may be overestimated. The reasons for the overestimation are detailed in 5.1.
@@ 5.1 Experimental Settings
Eval metrics. Followed by [2], we use the mean average precision at an IOU threshold of 0.5 ($A{P_{50}}$, %) as the metric. For PASCAL VOC2007, following previous works [17], we provide the $A{P_{50}}$ of the current task classes and the previous task classes to better reflect the method’s stability and plasticity. There are two evaluation methods for obtaining task precision: validating on the entire test set versus using task-specific test subsets. The second method yields higher precision for the same detector. We adopt the first method and add superscript to results from the second method to ensure fair comparison. For MS COCO, following previous works [16], we provide the $A{P_{50}}$ of all learned classes after learning at each task.
Implementation details. We implement our proposed method based on Deformable-DETR [47] pre-trained on the MS COCO dataset and Co-DETR [49] pre-trained on the Objects365 dataset [30], both obtained from HuggingFace. The large-scale Objects365 dataset contains 365 categories and over 600,000 images, making it a suitable pre-training source for incremental learning experiments on MS COCO dataset. Furthermore, since the official MD-DETR implementation is only available on Deformable-DETR, we port MD-DETR [2] on Co-DETR for a fair comparison.
@@ 5.2 Comparison
Table 4: Average precision ($A{P_{50}}$, %) is compared on the PASCAL VOC2007 dataset under multi-step settings of 10+5+5 and 5+5+5+5. We add the superscript to the accuracy may be overestimated. The reasons for the overestimation are detailed in 5.1.
@@
Single-step setting. We compare three single-step scenarios on the PASCAL VOC2007 dataset, where the co-occurrence levels gradually increase in the 19+1, 15+5, and 10+10 settings. As shown in Tab. 1, P2IOD with MS COCO and Objects365 pretrained detectors achieve excellent performance across all experimental settings. Compared to the prompt-pool-based MD-DETR, P2IOD achieves accuracy improvements of 1.6% / 0.8%, 2.5% / 3.9%, and 8.0% / 5.2% in the respective scenarios, indicating that the performance advantage of P2IOD becomes increasingly significant as the co-occurrence level rises. This trend further demonstrates that P2IOD mitigates the interference caused by prompts pool confusion in co-occurring scenarios. We also compare single-step scenarios on the MS COCO dataset in the Appendix B.1, and the results further demonstrate the effectiveness of P2IOD.
Multi-step setting. We compare the multi-step settings on PASCAL VOC2007 and MS COCO datasets. Tab. 4 shows that on PASCAL VOC2007, the performance degradation of MD-DETR becomes increasingly severe as the number of incremental steps grows. The degradation stems from prompts pool confusion which intensifies as the number of tasks increases. In contrast, our method effectively mitigates such confusion, consistently achieving superior performance across all settings. For the MS COCO dataset, as shown in Tab. 2, P2IOD also exhibits the aforementioned advantages. Moreover, P2IOD consistently outperforms other existing approaches across different settings on both datasets, demonstrating its effectiveness and the strong potential of prompt-based techniques in IOD.
@@ 5.3 Analysis
Ablation Study. In Tab. 3, categories 1-5 reflect the stability, while categories 6-20 mainly reflect plasticity. After introducing the pseudo-labeling method (b), due to the lack of learnable parameters, pseudo-labeling not only fails to enhance stability but also interferes with current task learning, reducing plasticity. The parameterized prompt structure (c) increases the accuracy of categories 6-20 by 9.5%, significantly enhancing plasticity, but the accuracy of categories 1-5 drops by 2.6%, indicating that forgetting still exists. The model fusion (d) alleviates the forgetting problem and balances stability and plasticity to some extent. The sparse loss (f) reduces parameter conflicts between tasks, improving overall accuracy by 1.1%. When combined with all methods, it increases the overall accuracy by 9.0% compared to the baseline. Furthermore, removing the pseudo-labeling (e) results in a significant decrease in stability. The observation demonstrates that our method, by alleviating the prompts pool confusion, allows the pseudo labeling mechanism to effectively mine old-class objects in the background without introducing adverse effects.
Impact of Hidden Layer Dimension. We analyze the impact of the hidden layer dimension in the parameterized prompt structure. The dimension of the hidden layer influences the degree of dimensionality reduction applied to the proposals. We conduct experiments in the PASCAL VOC2007 under the 5+5+5+5 setting. In Fig. 4, as the hidden layer dimension increases, the model’s accuracy initially improves and then declines, suggesting that a moderate increase in the hidden dimension helps retain critical information, while an overly large dimension introduces redundant information that interferes with prompt generation. As the hidden dimension increases, the number of parameters in the parameterized prompt increases accordingly. Our method achieves a favorable trade-off between performance and parameter efficiency (76.3%, 1.1M) when the hidden dimension is set to 64. In contrast, MD-DETR [2] requires more parameters, while simultaneously achieving lower accuracy (61.5%, 1.8M).
@@
Prompt Comparison. We compare the distribution similarity of prompts across tasks between P2IOD and MD-DETR [2]. We conduct this experiment on PASCAL VOC2007 and use Maximum Mean Discrepancy (MMD) [8] to evaluate the distribution similarity of prompts, with the average MMD (A-MMD) across all tasks as the evaluation metric. A larger A-MMD value indicates a more significant prompt diversity. As shown in Fig. 5, the diversity of prompts in our P2IOD increases with the depth of decoder layers, and the diversity at the final layer is significantly higher than that of MD-DETR. Our method generates category-independent prompts in shallow layers and category-related prompts in deep layers, aligning well with the multi-layer decoder architecture, while MD-DETR is constrained by its pool structure and struggles to match this characteristic. Furthermore, our method can generate more diverse prompts at the final layer used for object prediction. The variations in prompt distributions highlight the effectiveness of our approach.
@@ 5.4 Visualized Comparison
We analyze the visualized comparison between P2IOD and MD-DETR to illustrate that the confusion is being addressed. In Fig. 3, the visualizations of MD-DETR exhibit numerous false positives, indicating that the confused prompts pool introduces incorrect prompts into the detector, thereby increasing scores for irrelevant categories. In contrast, P2IOD reduces such false positives, demonstrating the effectiveness of our approach in mitigating confusion. We also observe that although P2IOD and MD-DETR have nearly identical confidence in detecting people in the first task, MD-DETR’s confidence rapidly declines as the number of tasks increases, suggesting that confusion in MD-DETR undermines confidence in positive objects. In contrast, the confidence in P2IOD remains almost unchanged, proving that our method is unaffected by confusion.
@@ 6 Conclusion
In this study, we identify a confusion issue within the prompt-pool-based IOD methods. To address this issue, we argue that prompts in IOD should not be exclusively assigned to individual tasks but should instead exhibit adaptive consolidation properties across tasks, with constrained updates. We propose a parameterized prompt structure and parameterized prompt fusion to validate our hypothesis. Experiments on multiple datasets demonstrate that our framework exhibits superior performance. To our knowledge, this is the first work addressing prompts pool confusion in IOD, laying a foundation for broader prompt-based IOD applications.
Acknowledgments
This work is partially supported by the Chinese Academy of Sciences Project for Young Scientists in Basic Research (YSBR-107).
@@ Acknowledgments
Appendix A Implementation Details
The pre-trained detectors use the original hyperparameter settings as those in their respective papers [47, 49]. We employ the Adam optimizer with a weight decay of 0.0001. The learning rate is set to 0.0001 for the class embeddings and the parametrized prompt structure, and a lower learning rate of 0.00001 is used for the bounding box embeddings, while freezing the remaining parameters. For training Deformable-DETR [47] on the PASCAL VOC2007 dataset [6], we train each incremental task for 100 epochs, with the learning rate dropped to 0.1 of its original value at the 80th epoch. For training Co-DETR [49] on the PASCAL VOC2007 dataset, we train each incremental task for 12 epochs, with the learning rate dropped to 0.1 of its original value at the 7th epoch. For training Co-DETR on the MS COCO dataset, we train each incremental task for 1 epoch. It is worth noting that in the 19+1 setting on PASCAL VOC2007, the second task contains only 279 images. Due to the limited data, Deformable-DETR tends to overfit during training. To mitigate this issue, we set the hyperparameters of the focal loss to $\alpha=0.5$ and $\gamma=3.0$. The reported accuracy is the average of three independent trials. All experiments are conducted on NVIDIA RTX 4090 GPUs.
@@ Appendix A Implementation Details
Appendix B Additional Experiment Results
B.1 Single-step Comparison on MS COCO dataset
We compare two single-step scenarios on the MS COCO dataset, namely the 40+40 and 70+10 settings. As shown in Tab. 5, P2IOD performs excellently across all experimental settings. Compared to the prompt-pool-based MD-DETR, P2IOD achieves $A{P_{50}}$ accuracy improvements of 5.4% and 6.3% in the two scenarios, demonstrating that the proposed method effectively addresses the prompts pool confusion.
B.2 Impact of Parameterized Prompt Fusion Threshold
@@
We quantitatively analyze the impact of the ${\text{top-}}k$ and ${\text{top-}}l$ in parameterized prompt fusion. The threshold determines the proportion of the previous and current parameterized prompt structures retained during fusion. To clarify the impact of one-step fusion on accuracy, we perform experiments in the PASCAL VOC2007 under 10+10 setting. As shown in Tab. 6, we observe that neither introducing model fusion (no fusion) nor using only the first task’s parameterized prompts (${\text{top-}}k=1.0$) results in poor performance. When ${\text{top-}}k=0.0$ and ${\text{top-}}l=0.0$, the method averages the consistent parameters, improving the performance compared to the non-fused approach, demonstrating that averaging consistent parameters enhances generalization. Furthermore, by comparing different values of ${\text{top-}}k$ and ${\text{top-}}l$, we find that moderately retaining parameters from both the current and previous tasks can further improve performance. The accuracy improvement indicates that retaining key parameters from each task helps preserve task-specific knowledge. We observe the best performance at ${\text{top-}}k=0.7$ and ${\text{top-}}l=0.3$, resulting in a 1.08% accuracy increase over the non-fused method.
B.3 Impact of $\lambda$ in Sparse Loss